XRCF 2ND FLOOR SHIFT REPORT Starting Date: 23 January 1997 Shift: B File Ref: /home/xrcfdude/shiftreports/970123.B RS/RE: D. Schwartz Operator: S. Vitek Software: J. Moran Shift Scientist: B. Wargelin Project Scientist: M. Weisskopf ==================== Measurement Summary: ------------------- 109987 D-IXS-EA-99.020 last DCM scan 109997 D-IXF-BG-1.026 Al-K background 109998 D-IXF-BG-1.026 Al-K background (with 244Cm in SSD_5) Restore old FOA for Al-K EIPS 110000 D-IXH-RF-1.005 (setup) beamcen fpc_x2 20-um Update FOA y and z values 110001 D-IXH-RF-1.005 (setup) beamcen ssd_x 70-um 110003 D-IXH-RF-1.005 (setup) beamcen ssd_x 20-um Updated master ssd_x aperture table--relative to fpc_x2 110004 D-IXH-RF-1.005 HSI Ring Focus 110008 D-IXF-FA-4.013 FPC Aging 110009 D-IXF-3D-66.001 FPC, Shells 13 110010 D-IXS-3D-66.009 SSD, Shells 13 110011 D-IXS-3D-66.010 SSD, Shells 46 110012 D-IXF-3D-66.002 FPC, Shells 46 110013 D-HXF-3D-66.004 FPC, Shells 46, HEG +1,+2,0,-1,-2 110014 D-HXS-3D-66.012 SSD, Shells 46, HEG +1,+2,0,-1,-2 [The following lines appeared in error so I've edited them - dd] 110015 D-HXF-3D-66.003 FPC, Shells 13, MEG +1,+2,+3,+4,0,-1,-2,-3,-4 110016 D-HXS-3D-66.011 SSD, Shells 13, MEG +1,+2,+3,+4,0,-1,-2,-3,-4 110017 D-LXS-3D-66.013 SSD, LETG, 0,+1,+3,-3,-1 110018 D-LXF-3D-66.005 FPC, LETG, 0,+1,+3,-3,-1 110019 D-LXS-3D-66.014 SSD, LETG, + and - 2 to 7 Results: ------- The FPC Aging test, which was done correctly for the first time, showed that there is no observable degradation of fpc_x2 performance. Supplementary Tests: ------------------- Beamcens with fpc_x2/20-um, ssd_x/70-um, ssd_x/20-um were done as part of the setup for the Ring Focus test. The results were used to update the FOA and ssd_x aperture table. There was no change in z, but 217 um were added to all ssd_x y positions. This change was required because of the change in ssd aperture-wheel direction-of- approach made last week. Notes: ----- The Project Science Measurement Summaries do not list grating orders in chronological order (in the little picture of the OTGs), although the 2-axis projection plots do show the order correctly. The D-xSx-3D-26.xxx series was changed to -66.xxx because the -26.xxx series had been run previously. The D-xXx-3D-66.xxx tests which used very high fluxes and only high orders had to be modified to avoid moving the FPC and SSD through 0 order. The FPC Aging test had to be modified to use fpc_x2; the CMDB listed fpc_x1. Problems: -------- Midway through this shift's Al-K tests, a narrow low-energy peak was observed in the fpc_5 spectra. Looking back through previous runs, the problem appeared when the source was changed from the DCM to EIPS Al-K. During C-shift many tests were conducted to try and understand this problem. Observations include: the "noise" rate is proportional to aperture size, it only appears when the fpc is exposed to x-rays, the noise appears WHETHER THE DETECTOR HIGH VOLTAGE IS ON OR OFF, the noise peak moves to higher channels as gas pressure is reduced, and scanning the 1-mm aperture along the anode wire has no effect. Judging from the pulser signal, the amplifier is working fine with the correct gain, rebooting the PC server has no effect, dis/reconnecting all signal and power cables has no effect, no does the SSD HV have any effect. More x-ray tests were conducted during C-shift... XRCF 2ND FLOOR SHIFT REPORT Starting Date: 23 January 1997 Shift: C File Ref: /home/xrcfdude/shiftreports/970123.C RS/RE: EMK Operator: MO Software: Shift Scientist: MS Project Scientist: SO ==================== Measurement Summary: ------------------- 110020 D-LXS-3D-66.015 SSD, 3D Scan LETG +,- 7 to 14 @ Al-K 110021 D-LXS-3D-66.016 SSD, 3D Scan LETG +,- 15 to 25 @ Al-K 110022 D-LXF-3D-66.006 FPC2, 3D Scan LETG +,- 3 to 7 @ Al-K 110023 D-LXF-3D-66.007 FPC2, 3D Scan LETG +,- 7 to 14 @ Al-K 110024 D-LXF-3D-66.008 FPC2, 3D Scan LETG +,- 15 to 25 @ Al-K 110036 D-IXF-BG-1.021 detector backgrounds for Be-K measurments ------ D-IXF-BU-2.041 Be-K beam uniformity - setup aborted due to fpc_5 issues 110063 D-IXF-BG-1.010 detector backgrounds for Sn-La 110064 D-BND-BU-2.020 Sn-La beam uniformity Supplementary Tests: ------------------- Many diagnostic tests were done to determine whether the fpc_5 problems were caused by (inductive?) coupling with EIPS, detector electronics or gain. Detector gain appears to be declining. Problems: -------- The additional tests that we performed on fpc_5 indicate that the "noise" observed during the Be measurements was the Be spectrum, and that what had happened was that the fpc_5 gain had degraded from N(E peak channel)/sqrt[E(keV)] = 79 to now about 24-25. This degradation was apparent in the later DCM scans, though not as bad (more like 50). We did not search through the data to determine when this started. We should reverify the claim from the earlier shift that this "noise" on fpc_5 was present with or without voltage on fpc_5.